OpenStudy (anonymous):

are the three numbers 72,68,12 measurements of a right triangle

OpenStudy (anonymous):

your job is to see whether \[12^2+68^2=72^2\] i would use a calculator

OpenStudy (anonymous):

144 64 5184

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

not that you could also divide everything by 4 and check to see if \(4,17,18\) were sides of a right triangle the answer is no because \[4^2+17^2=305\neq 18^2\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw